# Exotic synchronization in continuous time crystals outside the symmetric subspace

### v1

@@ Abstract
Exploring continuous time crystals (CTCs) within the symmetric subspace of spin systems has been a subject of intensive research in recent times. Thus far, the stability of the time-crystal phase outside the symmetric subspace in such spin systems has gone largely unexplored. Here, we investigate the effect of including the asymmetric subspaces on the dynamics of CTCs in a driven dissipative spin model. This results in multistability, and the dynamics becomes dependent on the initial state. Remarkably, this multistability leads to exotic synchronization regimes such as chimera states and cluster synchronization in an ensemble of coupled identical CTCs.
Introduction.—
@@
Crystals have traditionally been characterized by their spatial periodicity, wherein atoms are arranged in a repeating pattern. Time crystals showcase a distinct form of periodicity, namely in the temporal dimension Wilczek (2012); Watanabe and Oshikawa (2015); Zaletel et al. (2023); Sacha and Zakrzewski (2017). Initially, time crystals breaking discrete time-translation symmetry were conceptualized and termed discrete time crystals Sacha (2015); Else et al. (2016); Khemani et al. (2016); Russomanno et al. (2017); Surace et al. (2019); Heugel et al. (2019); Khasseh et al. (2019); Sakurai et al. (2021); Sacha and Zakrzewski (2017); Zaletel et al. (2023). A more recent development in the field has been the emergence of continuous time crystals (CTCs) Iemini et al. (2018); Tucker et al. (2018); Buča et al. (2019); Booker et al. (2020); Buča and Jaksch (2019); Zhu et al. (2019); Lledó et al. (2019); Seibold et al. (2020); Prazeres et al. (2021); Piccitto et al. (2021); Carollo and Lesanovsky (2022); Hajdušek et al. (2022); Krishna et al. (2023a); Hurtado-Gutiérrez et al. (2020); Cosme et al. (2023); Buonaiuto et al. (2021); Cabot et al. (2023a), which break continuous time-translation symmetry. CTCs were initially discussed in the context of a driven dissipative Dicke model Iemini et al. (2018). Their behavior was also studied in other spin systems like collective $d$-level systems Prazeres et al. (2021), $p$-$q$ interacting spin models Piccitto et al. (2021), spin star models Krishna et al. (2023a), and spin systems with infinite-range interactions Cosme et al. (2023). In addition to the time-crystal behavior, the thermodynamics Carollo et al. (2023) of driven Dicke models and the occurrence of entanglement phase transitions Passarelli et al. (2022); Cabot et al. (2023b) have also been investigated recently. The application of time-crystal phases in sensing Cabot et al. (2023c); Montenegro et al. (2023) and quantum engines Carollo et al. (2020) has been proposed. Furthermore, experimental platforms such as Bose-Einstein condensate systems have been used to realize CTCsKeßler et al. (2021); Kongkhambut et al. (2022). All of these examples are variants of the Dicke model: the system is a collective spin that is built up of $N$ spins 1/2. These systems are characterized by permutational symmetry which is a strong symmetry of the underlying open quantum system Buča and Prosen (2012); Albert and Jiang (2014a). Previous investigations of both superradiance Dicke (1954); Andreev et al. (1980); Wang and Hioe (1973); Scully and Svidzinsky (2009) and time-crystal behavior have focused on the totally symmetric Dicke subspace, which corresponds to the maximum-excitation subspace. The assumption of permutational symmetry is not always straightforward to enforce under realistic conditions Ferioli et al. (2023). This leads to the important open question of the role played by the lower-excitation subspaces involving asymmetric Dicke states in the dynamics of time crystals. Recent studies have shown that asymmetric subspaces are crucial for understanding subradiant phases Bhatti et al. (2018); Bojer and von Zanthier (2022); Gegg et al. (2018); Wiegner et al. (2011) and the thermodynamics of permutation-invariant quantum many-body systems Latune et al. (2019, 2020); Yadin et al. (2023). In this Letter, we investigate the existence of time crystals outside the symmetric Dicke subspace in a spin-only description of the driven dissipative Dicke model Iemini et al. (2018); Hajdušek et al. (2022). We show that incorporating the asymmetric subspaces enhances the complexity and richness of CTC dynamics and yields an extended phase diagram. Using mean-field analysis, we find that CTCs exhibit multistability in the asymmetric subspaces, which was not observed for CTCs restricted to the symmetric subspace. The existence of multistability is confirmed by analyzing the eigenspectra of the Liouville superoperator that describes the Markovian dynamics of open quantum systems Buča and Prosen (2012); Albert and Jiang (2014b); Krishna et al. (2023b). Depending on the system parameters, an initial state outside the symmetric subspace can always be found which breaks time-translational symmetry. Interestingly, this initial-state dependence leads to exotic synchronization phenomena like chimera states in a network of coupled identical CTCs that are characterized by the coexistence of synchronized and unsynchronized regions Kuramoto and Battogtokh (2002); Abrams and Strogatz (2004). Furthermore, such a network of CTCs also exhibits cluster synchronization, where identical oscillators arrange themselves into two or more differently synchronized clusters Haugland (2021).
@@
Figure 1: (a) Phase diagram of a continuous time crystal (CTC) characterized using fixed-point analysis of mean-field equations with $m=\langle\hat{S}\rangle/S$. (b) Phase-space portrait of the CTC for various initial states (indicated by different symbols in (a)) for $\Omega/\kappa=0.75$. (c) Schematic representation of block diagonal structure of Liouville superoperator $\mathcal{L}=\mathcal{L}_{S}+\mathcal{L}_{A}$ where $\mathcal{L}_{S(A)}$ corresponds to symmetric (asymmetric) Dicke subspace. Here, dotted lines separate the blocks with $m>\Omega/\kappa$ from those with $m<\Omega/\kappa$. Only the sub-blocks having rescaled magnetization value $m<\Omega/\kappa$ will have purely imaginary eigenvalues in the thermodynamic limit $S\rightarrow\infty$. (d) Oscillations of $m_{z}$ for an initial state prepared in the asymmetric subspace $m=0.1$ for $\Omega/\kappa=0.9$.
@@
We begin with an investigation of the mean-field dynamics of a collective spin using fixed-point analysis, which describes the steady-state dynamics of the system. Then, we will use Liouville superoperator theory to validate the mean-field results. Finally, we discuss the existence of a chimera state and cluster synchronization in coupled CTCs before summarizing our results.
Mean-field analysis and phase diagram.—
@@
The master equation governing the dynamics of the CTC Iemini et al. (2018); Hajdušek et al. (2022) is given by
@@
where $\mathcal{L}_{\Omega,\kappa}$ is the Liouville superoperator, $\hat{S}_{\alpha}=\sum_{i}\sigma^{i}_{\alpha}/2$ are collective spin operators where $\alpha\in\{x,y,z\}$, $S=N/2$ is the maximum spin value for $N$ spins $1/2$ and $\hat{S}_{\pm}=\hat{S}_{x}\pm i\hat{S}_{y}$. The dynamics of the given system in the thermodynamic limit ($N\rightarrow\infty$) can be well understood using mean-field analysis. The evolution of mean-field expectation values of the rescaled spin operators, $m_{\alpha}=\langle\hat{S}_{\alpha}\rangle/S$ is governed by
@@
The total spin of the system is conserved since $\hat{S}^{2}$ commutes with all other spin operators in Eq. (1) and hence is a strong symmetry of the system Buča and Prosen (2012); Albert and Jiang (2014b); Manzano and Hurtado (2018). At the mean-field level, this implies that $m^{2}_{x}+m^{2}_{y}+m^{2}_{z}=m^{2}$ is a conserved quantity since correlations are assumed to vanish. For the symmetric Dicke subspace, $m$ is constrained to take the value $1$, which led to the breaking of time-translational symmetry at $\Omega/\kappa=1$ in earlier studies Iemini et al. (2018); Hajdušek et al. (2022). We will now consider the more general case $10\leq m<$ 1. The fixed points, $(m_{x}^{*},m_{y}^{*},m_{z}^{*})$, of the coupled differential Eqs. (Mean-field analysis and phase diagram.—) for this case are given as $\mathcal{M}_{1}=(\pm m\sqrt{1-(m\kappa/\Omega)^{2}},m^{2}\kappa/\Omega,0)$ and $\mathcal{M}_{2}=(0,\Omega/\kappa,\pm\sqrt{m^{2}-(\Omega/\kappa)^{2}})$. The stability of these fixed points is given by the eigenvalues of the Jacobian $\mathcal{J}$, whose matrix elements are defined as $\mathcal{J}_{\alpha\beta}=d\dot{m}_{\alpha}/dm_{\beta}$. The eigenvalues of $\mathcal{J}$ corresponding to fixed points $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\{\Lambda_{1}\}=\{0,\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}}\}$ and $\{\Lambda_{2}\}=\{0,\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}},\pm\kappa\sqrt{m^{2}-(\Omega/\kappa)^{2}}\}$ respectively, where both the upper or both the lower signs have to be chosen together. Note that the dynamics of the system are now dependent on both the system parameters $\Omega/\kappa$ and rescaled total magnetization $m$ of the given spin-subspace. For $\Omega/\kappa<m$, $\mathcal{M}_{2}$ is the only physical state of the system since $\mathcal{M}_{1}$ has unphysical imaginary values for the expectation of spin components. The corresponding eigenvalues $\{\Lambda_{2}\}$ for the fixed point $\mathcal{M}_{2}$ are purely real, and hence it is a saddle fixed point of the system. Therefore, the system relaxes to a time-invariant state given by the fixed point $\mathcal{M}_{2}$ for initial states having $m>\Omega/\kappa$, and no time-crystal behavior is observed. For states with $m<\Omega/\kappa$, $\mathcal{M}_{1}$ is the only physical fixed point of the system. The eigenvalues $\{\Lambda_{1}\}$ for the fixed point $\mathcal{M}_{1}$ are purely imaginary, and hence, the corresponding fixed point is a center. Therefore, the system oscillates in time for initial states having $m<\Omega/\kappa$, and time-translational symmetry is broken even for $\Omega/\kappa<1$. Using this fixed-point analysis, we can characterize the phase transition of the system as described by Fig. 1(a). The system exhibits multistability for $\Omega/\kappa<1$. The two fixed points $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have different basins of attraction: depending upon the initial state, the system can exhibit two different phases, namely, a melted and a time-crystal phase. For the Dicke subspace, this bistability does not exist, as depicted in Fig. 1(a). The Dicke subspace corresponds to initial states having $m=1$, which shows that for $\Omega/\kappa<1$, the system always settles down to a saddle fixed point (dashed red line), and for $\Omega/\kappa>1$, it shows oscillatory behavior (solid red line). However, for $m<1$, both time-translation invariant and symmetry-broken phases coexist and are separated by the blue line corresponding to $m=\Omega/\kappa$. Initial states having $m>\Omega/\kappa$ on the left of $m=\Omega/\kappa$ (blue) line settle down to the saddle fixed point $\mathcal{M}_{2}$. The parameter region right to the $m=\Omega/\kappa$ (blue) line corresponds to $m<\Omega/\kappa$ and gives stable oscillations around the fixed point $\mathcal{M}_{1}$ leading to time-translational symmetry breaking. To illustrate the phase-space portrait of the fixed points discussed above, we consider the following coordinate transformation $P=\tan^{-1}\left[m_{y}/m_{x}\right],~{}Q=m_{z}$ Iemini et al. (2018). In terms of the new co-ordinates $P$ and $Q$, the fixed points $\{P^{*},Q^{*}\}$ are given as $\mathcal{M}^{P,Q}_{1}=(\tan^{-1}[\pm 1/\sqrt{(\Omega/m\kappa)^{2}-1}],0)$ and $\mathcal{M}^{P,Q}_{2}=(\pm\pi/2,\pm\sqrt{m^{2}-(\Omega/\kappa)^{2}})$. The phase-space portrait of the CTC for different $m$ values is depicted in Fig. 1(b) for $\Omega/\kappa=0.75$. Initial states having $m<\Omega/\kappa$ keep oscillating in the vicinity of the fixed point $\mathcal{M}^{P,Q}_{1}$. As soon as we choose an initial state with $m>\Omega/\kappa$, the corresponding trajectory in Fig. 1(b) settles down to the saddle fixed point $\mathcal{M}^{P,Q}_{2}$. For $\Omega/\kappa>1$, there exists no physical state of the system that can have $m>\Omega/\kappa$ since $m\leq 1$ for the rescaled spin operators. The stability of the oscillations around the fixed point $\mathcal{M}_{1}^{P,Q}$ can be verified by observing Fig. 1(b), where each limit cycle trajectory consists of around 4000 cycles of oscillations. These trajectories are obtained by the time evolution of the exact mean-field equations (Mean-field analysis and phase diagram.—) (including non-linear terms) and show no deviation even at such long times.
Role of asymmetric subspaces.—
@@
We now go beyond mean-field theory and study the full master equation Eq. (1). The Hilbert space of the ensemble consisting of $N$ spins can be described as a direct sum of irreducible representations of $SU(2)$ using the Clebsch-Gordan decomposition Sakurai and Commins (1995). Since $\hat{S}^{2}$ is a strong symmetry of the dynamics, the Liouville superoperator block-diagonalizes Buča and Prosen (2012); Thingna and Manzano (2021) in respective spin-$J$ eigenbases such that $\mathcal{L}=\oplus\mathcal{L}_{J}$, as shown in Fig. 1(c). Therefore, the dynamics is constrained within each spin-$J$ subspace (invariant subspaces of dynamics). The maximum angular momentum value $J=S=N/2$ corresponds to the totally symmetric Dicke subspace; lower $J$ values correspond to asymmetric subspaces. The spin-$J$ irreducible representation has dimension $12J+$ 1 and multiplicity $n_{J}=N!(2J+1)/(N/2+J+1)!(N/2-J)!$ Georgi (2000); Ramadevi and Dubey (2019); Yadin et al. (2023). The total Hilbert space is spanned by the generalized Dicke basis, $\{|J,m_{J},\mu_{J}\rangle\}$. Here, $J\in\{S,S-1\ldots J_{0}\}$, $m_{J}\in\{J,J-1\ldots-(J-1),-J\}$ and $1\leq\mu_{J}\leq n_{J}$, with $J_{0}=0\text{ or }1$ depending on whether $N$ is even or odd. We can associate each of the spin-$J$ subspaces with the mean-field total angular momentum value $m=J/S$. The states with $m<1$ belong to the asymmetric angular momentum sectors.
@@
The mean-field analysis suggests that for any given value of $\Omega/\kappa$, we can always find an initial state having $m<\Omega/\kappa$ such that the system oscillates in time. In terms of parameters rescaled by $m$, the critical point is located at $\Omega/(m\kappa)=1$ for each invariant subspace. The well-studied case with $J=S$ describing the symmetric subspace (with dimension $N+1$) has $m=1$ and a critical point at $\Omega/\kappa=1$. The remainder of the Hilbert space with dimension $2^{N}-(N+1)$ spans the asymmetric subspaces and has $m<1$. Hence, even for $\Omega/\kappa<1$, the exact dynamics of the finite-size system exhibits time-crystal behavior. Figure 1(d) depicts the time evolution with the initial state belonging to the $m=0.1$ subspace, having $\{m_{x},m_{y},m_{z}\}=\{0,0,0.1\}$ for $\Omega/\kappa=0.9$. As the system size increases, the oscillation in $m_{z}$ becomes persistent and approaches the mean-field limit. Any initial state belonging to the $m>0.9$ subspace will decay and go to a time-invariant steady state irrespective of system size. Thus, the exact treatment confirms the mean-field analysis for $\Omega/\kappa<1$.
Figure 3: Normalized Fourier transform $\mathcal{F}[m^{z}_{\alpha}]$ (a-c) and time evolution of $m^{z}_{\alpha}$ (d-f) for $n=20$ continuous time crystals (CTCs). All the CTCs are identical ($\Omega_{\alpha}/\kappa=0.9$ and $\kappa=1$) and are initialized in different $m$ subspaces. The two ensembles $\varepsilon_{1,2}$ are represented by the colors black and red above subfigures (a,b,c), and the CTCs within one ensemble by different intensities. Subfigures (a,d) show the case of uncoupled time crystals. Coupling CTCs within each ensemble with a coupling strength of $\Gamma/\kappa=0.35$ results in the synchronization of CTCs within the ensemble as shown in (b,e). In subfigures (c,f) all CTCs are coupled. Surprisingly, CTCs in ensemble $\varepsilon_{1}$ get synchronized, while the CTCs in the ensemble $\varepsilon_{2}$ remain unsynchronized: the system exhibits a chimera state.
@@
Chimera state in coupled CTCs.—
We now discuss the effect of including asymmetric subspaces on the dynamics of $n$ coherently coupled CTCs described by the following master equation,

### v2

@@ Abstract
Exploring continuous time crystals (CTCs) within the symmetric subspace of spin systems has been a subject of intensive research in recent times. Thus far, the stability of the time-crystal phase outside the symmetric subspace in such spin systems has gone largely unexplored. Here, we investigate the effect of including the asymmetric subspaces on the dynamics of CTCs in a driven dissipative spin model. This results in multistability, and the dynamics becomes dependent on the initial state. Remarkably, this multistability leads to exotic synchronization regimes such as chimera states and cluster synchronization in an ensemble of coupled identical CTCs. Interestingly, it leads to other nonlinear phenomena such as oscillation death and signature of chaos.
@@
Introduction.— Crystals have traditionally been characterized by their spatial periodicity, wherein atoms are arranged in a repeating pattern. Time crystals showcase a distinct form of periodicity, namely in the temporal dimension Wilczek (2012); Watanabe and Oshikawa (2015); Zaletel et al. (2023); Sacha and Zakrzewski (2017). Initially, time crystals breaking discrete time-translation symmetry were conceptualized and termed discrete time crystals Sacha (2015); Else et al. (2016); Khemani et al. (2016); Russomanno et al. (2017); Surace et al. (2019); Heugel et al. (2019); Khasseh et al. (2019); Sakurai et al. (2021); Sacha and Zakrzewski (2017); Zaletel et al. (2023). A more recent development in the field has been the emergence of continuous time crystals (CTCs) Iemini et al. (2018); Tucker et al. (2018); Buča et al. (2019); Booker et al. (2020); Buča and Jaksch (2019); Zhu et al. (2019); Lledó et al. (2019); Seibold et al. (2020); Prazeres et al. (2021); Piccitto et al. (2021); Carollo and Lesanovsky (2022); Hajdušek et al. (2022); Krishna et al. (2023a); Hurtado-Gutiérrez et al. (2020); Cosme et al. (2023); Buonaiuto et al. (2021); Cabot et al. (2023a), which break continuous time-translation symmetry. CTCs were initially discussed in the context of a driven dissipative Dicke model Iemini et al. (2018). Their behavior was also studied in other spin systems like collective $d$-level systems Prazeres et al. (2021), $p$-$q$ interacting spin models Piccitto et al. (2021), spin star models Krishna et al. (2023a), and spin systems with infinite-range interactions Cosme et al. (2023). In addition to the time-crystal behavior, the thermodynamics Carollo et al. (2024) of driven Dicke models and the occurrence of entanglement phase transitions Passarelli et al. (2022); Cabot et al. (2023b) have also been investigated recently. The application of time-crystal phases in sensing Cabot et al. (2024); Montenegro et al. (2023) and quantum engines Carollo et al. (2020) has been proposed. Furthermore, experimental platforms such as Bose-Einstein condensate systems have been used to realize CTCs Keßler et al. (2021); Kongkhambut et al. (2022).
All of these examples are variants of the Dicke model: the system is a collective spin that is built up of $N$ spins 1/2. These systems are characterized by permutational symmetry which is a strong symmetry of the underlying open quantum system Buča and Prosen (2012); Albert and Jiang (2014a). Previous investigations of both superradiance Dicke (1954); Andreev et al. (1980); Wang and Hioe (1973); Scully and Svidzinsky (2009) and time-crystal behavior have focused on the totally symmetric Dicke subspace, which corresponds to the maximum-excitation subspace. The assumption of permutational symmetry is not always straightforward to enforce under realistic conditions Ferioli et al. (2023). This leads to the important open question of the role played by the lower-excitation subspaces involving asymmetric Dicke states in the dynamics of time crystals. Recent studies have shown that asymmetric subspaces are crucial for understanding subradiant phases Bhatti et al. (2018); Bojer and von Zanthier (2022); Gegg et al. (2018); Wiegner et al. (2011) and the thermodynamics of permutation-invariant quantum many-body systems Latune et al. (2019, 2020); Yadin et al. (2023).
In this Letter, we investigate the existence of time crystals outside the symmetric Dicke subspace in a spin-only description of the driven dissipative Dicke model Iemini et al. (2018); Hajdušek et al. (2022). We show that incorporating the asymmetric subspaces enhances the complexity and richness of CTC dynamics and yields an extended phase diagram. Using mean-field analysis, we find that CTCs exhibit multistability in the asymmetric subspaces, which was not observed for CTCs restricted to the symmetric subspace. The existence of multistability is confirmed by analyzing the eigenspectra of the Liouville superoperator that describes the Markovian dynamics of open quantum systems Buča and Prosen (2012); Albert and Jiang (2014b); Krishna et al. (2023b). Depending on the system parameters, an initial state outside the symmetric subspace can always be found which breaks time-translational symmetry. Interestingly, this initial-state dependence leads to exotic synchronization phenomena in a network of coupled identical CTCs like chimera states that are characterized by the coexistence of synchronized and unsynchronized regions Kuramoto and Battogtokh (2002); Abrams and Strogatz (2004); Yeldesbay et al. (2014); Chandrasekar et al. (2015); Ujjwal et al. (2017a). Furthermore, such a network of CTCs also exhibits cluster synchronization, where identical oscillators arrange themselves into two or more differently synchronized clusters Haugland (2021). Finally, we show that this highly nonlinear coupled system also features chaotic mean-field dynamics and the phenomenon of oscillation death.
@@
Figure 1: (a) Phase diagram of a continuous time crystal (CTC) characterized using fixed-point analysis of mean-field equations with $m=\langle\hat{S}\rangle/S$. (b) Schematic representation of block diagonal structure of Liouville superoperator $\mathcal{L}=\mathcal{L}_{S}+\mathcal{L}_{A}$ where $\mathcal{L}_{S(A)}$ corresponds to symmetric (asymmetric) Dicke subspace. Here, dotted lines separate the blocks with $m>\Omega/\kappa$ from those with $m<\Omega/\kappa$. Only the sub-blocks having rescaled magnetization value $m<\Omega/\kappa$ will have purely imaginary eigenvalues in the thermodynamic limit $N=2S\rightarrow\infty$.
@@
We begin with an investigation of the mean-field dynamics of a collective spin using fixed-point analysis, which describes the steady-state dynamics of the system. Then, we will use Liouville superoperator theory to validate the mean-field results. Finally, we discuss the existence of a chimera state, cluster synchronization, oscillation death and chaos in coupled CTCs before summarizing our results.
@@
Mean-field analysis and phase diagram.— The master equation governing the dynamics of the CTC Iemini et al. (2018); Hajdušek et al. (2022) is given by
@@
where $\mathcal{L}_{\Omega,\kappa}$ is the Liouville superoperator, $\hat{S}_{\alpha}=\sum_{i}\sigma^{i}_{\alpha}/2$ are collective spin operators where $\alpha\in\{x,y,z\}$, $S=N/2$ is the maximum spin value for $N$ spins $1/2$ and $\hat{S}_{\pm}=\hat{S}_{x}\pm i\hat{S}_{y}$. The dynamics of the system in the thermodynamic limit ($N\rightarrow\infty$) can be well understood using mean-field analysis. The evolution of mean-field expectation values of the rescaled spin operators, $m_{\alpha}=\langle\hat{S}_{\alpha}\rangle/S$ is governed by
@@
The total spin of the system is conserved since $[\hat{S}^{2},S_{x,y,z,\pm}]=0$ and hence is a strong symmetry of the system Buča and Prosen (2012); Albert and Jiang (2014b); Manzano and Hurtado (2018). At the mean-field level, this implies that $m^{2}_{x}+m^{2}_{y}+m^{2}_{z}=m^{2}$ is a conserved quantity since correlations are assumed to vanish Mukherjee et al. (2024).
Using the fixed-point analysis, we can characterize the phase transition of the system as described by Fig. 1(a) (see sup for details). Initial states on the left of the $m=\Omega/\kappa$ (blue) line ($m>\Omega/\kappa$) settle down to a saddle fixed point. In the parameter region $m<\Omega/\kappa$, the systems exhibit stable oscillations around the center fixed point leading to time-translational symmetry breaking. The frequency of oscillations is determined by the imaginary eigenvalues of the Jacobian matrix ±$\pm i\kappa\sqrt{(\Omega/\kappa)^{2}-m^{2}}$. The symmetric Dicke subspace corresponds to $m=1$, which was shown to lead to the breaking of time-translational symmetry at $\Omega/\kappa=1$ in earlier studies Iemini et al. (2018); Hajdušek et al. (2022). The system exhibits multistability for $\Omega/\kappa<1$. Depending upon the initial state, the system can exhibit two different phases, namely, a melted and a time-crystal phase.
@@
Role of asymmetric subspaces.— We now go beyond mean-field theory and study the full master equation Eq. (1). The Hilbert space of the ensemble consisting of $N$ spins can be described as a direct sum of irreducible representations of $SU(2)$ using the Clebsch-Gordan decomposition Sakurai and Commins (1995). Since $\hat{S}^{2}$ is a strong symmetry of the dynamics, the Liouville superoperator block-diagonalizes Buča and Prosen (2012); Thingna and Manzano (2021) in respective spin-$J$ eigenbases such that $\mathcal{L}=\oplus\mathcal{L}_{J}$, as shown in Fig. 1(b). Therefore, the dynamics is constrained within each spin-$J$ subspace (invariant subspaces of dynamics). The maximum angular momentum value $J=S=N/2$ corresponds to the totally symmetric Dicke subspace; lower $J$ values correspond to asymmetric subspaces. The spin-$J$ irreducible representation has dimension $12J+$ 1 and multiplicity $n_{J}=N!(2J+1)/(N/2+J+1)!(N/2-J)!$ Georgi (2000); Ramadevi and Dubey (2019); Yadin et al. (2023). The total Hilbert space is spanned by the generalized Dicke basis, $\{|J,m_{J},\mu_{J}\rangle\}$. Here, $J\in\{S,S-1\ldots J_{0}\}$, $m_{J}\in\{J,J-1\ldots-(J-1),-J\}$ and $1\leq\mu_{J}\leq n_{J}$, with $J_{0}=0\text{ or }1$ depending on whether $N$ is even or odd. We can associate each of the spin-$J$ subspaces with the mean-field total angular momentum value $m=J/S$. The states with $m<1$ belong to the asymmetric angular momentum sectors.
@@
Figure 3: Normalized Fourier transform $\mathcal{F}[m^{z}_{\alpha}]$ (a-c), the time evolution of $m^{z}_{\alpha}$ (d-f) and Pearson’s correlation matrix (g-i) for $n=20$ continuous time crystals (CTCs). All the CTCs are identical ($\Omega_{\alpha}/\kappa=0.9$ and $\kappa=1$) and are initialized in different $m$ subspaces. The two ensembles $\varepsilon_{1,2}$ are represented by the colors black ($n_{\varepsilon_{1}}\coloneqq\{1,2,\ldots,10\}$) and red ($n_{\varepsilon_{2}}\coloneqq\{11,12,\ldots,20\}$) above subfigures (a,b,c), and the CTCs within one ensemble by different intensities. The frequencies of the CTCs are uniformly sampled in the window from $0.2$ to $0.25$ in ensemble $\varepsilon_{1}$ and from $0.75$ to $0.85$ in ensemble $\varepsilon_{2}$. Subfigures (a,d,g) show the case of uncoupled time crystals. Coupling CTCs within each ensemble with a coupling strength of $\Gamma/\kappa=0.35$ results in the synchronization of CTCs within the ensemble, as shown in (b,e,h). In subfigures (c,f,i), all CTCs are coupled. Surprisingly, CTCs in ensemble $\varepsilon_{1}$ get synchronized, while the CTCs in the ensemble $\varepsilon_{2}$ remain unsynchronized: the system exhibits a chimera state.
@@
Figure 4: Subfigures (a-d) represent the cluster synchronization where (a,c) describe the dynamics of uncoupled CTCs ($\Gamma=0$), which get synchronized in different clusters for $\Gamma=0.35$ as shown in (b,d). The frequencies of the uncoupled CTCs in (a) have a uniform distribution from $0.2$ to $0.25$ in ensemble $\varepsilon_{1}$ and from $0.75$ to $0.8$ in ensemble $\varepsilon_{2}$. Various exotic synchronization regimes are characterized by investigating (e) the average value of the Pearson coefficient and (f) the maximum Lyapunov exponent. They are studied as a function of coupling strength $\Gamma$ and detuning $\Delta$ between the mean frequencies of two Gaussian-distributed ensembles $\varepsilon_{1,2}$, for $n=100$ CTCs with standard deviations $\sigma_{1}=\sigma_{2}=0.1$. Here $\Omega_{\alpha}/\kappa=0.9$ and $\kappa=1$ for all subfigures.
Chimera state and cluster synchronization.— We now discuss the effect of including asymmetric subspaces on the dynamics of $n$ coherently coupled CTCs described by the following master equation,